The temperature of an ideal gas is doubled from 200 K to 400 K while the volume is kept constant. Aaron was smart and looked up the kinetic theory. Which changes occur to the particles that make up the gas?
1
Expert's answer
2020-04-18T07:07:13-0400

According to a combined gas law:

p1 × V1 / T1 = p1 × V2 / T2

where p - pressure, V - volume , T - temperature.

As the volume is kept constant:

p1 / T1 = p1 / T2

If T1 = 200 K and T2 = 400 K:

p1 / 200 = p2 / 400

From here:

p1 = p2 / 2

p2 = 2p1

As a result, as the temperature of the ideal gas is doubled and the volume is kept constant, the pressure of the gas is also doubled. The process is caused by the increased kinetic energy of particles that move wit higher speed.